OpenStudy (anonymous):

eliminate the parameter and write the corresponding rectangular equation whose graph represents the curve. x=e^2t Y=e^2t

OpenStudy (jamesj):

So you're looking for an equation of the form: y = f(x), a function just in x. Looking at your equations, can you see what y is equal to?

OpenStudy (anonymous):

y=x right?

OpenStudy (jamesj):

Exactly

OpenStudy (anonymous):

crap I just realized I wrote it wrong x=e^2t and y =e^t sorry X0

OpenStudy (jamesj):

Well, that's still not so bad.

OpenStudy (jamesj):

what do you think?

OpenStudy (jamesj):

Remember that e^(2t) = e^t times e^t

OpenStudy (anonymous):

y=sqrt(x)?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

awesome thanks :)